# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ D.1 Drift kernels
The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)$; for trader $k$:
$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
@@ D.2 Deviation propagation
$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
@@ D.3 First-order condition
The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$). The conditional future demand in Volterra form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}(\tau,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes). Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3). $D^{i}(t,u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$. To pin down $D^{i}(t,u)$, we differentiate both sides in $t$.
@@ D.4 Leibniz differentiation and backward ODE
$\frac{d}{dt}H^{i,\mathrm{imp}}(t,u)=-\lambda(t)\,D^{i}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (D.7)
where the boundary term $\delta P_{t}\cdot D^{i}(t,u)=-\lambda(t)D^{i}(t,u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{i}_{DD}$ and where $D^{i}(t,u)$ first appears explicitly.
@@
In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}(t,u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$. Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
$\frac{d}{dt}H^{i,\mathrm{val}}(t,u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
@@
Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}(t,u)$ yields (5.4). The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$. Convexity (from $\lambda(t)>0$) ensures sufficiency.

### 18 Mar 2026 v2

@@ D.1 Drift kernels
The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)$; for trader $k$:
$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
@@ D.2 Deviation propagation
$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
@@ D.3 First-order condition
The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$). The conditional future demand in noise-state linear form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}_{\tau}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes). Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3). $D^{i}_{t}(u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$. To pin down $D^{i}_{t}(u)$, we differentiate both sides in $t$.
@@ D.4 Leibniz differentiation and backward ODE
$\frac{d}{dt}H^{i,\mathrm{imp}}_{t}(u)=-\lambda(t)\,D^{i}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (D.7)
where the boundary term $\delta P_{t}\cdot D^{i}_{t}(u)=-\lambda(t)D^{i}_{t}(u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{DD,i}$ and where $D^{i}_{t}(u)$ first appears explicitly.
@@
In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}_{t}(u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$. Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
$\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
@@
Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}_{t}(u)$ yields (5.4). The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$. Convexity (from $\lambda(t)>0$) ensures sufficiency.